OpenStudy (rmrjr22):

r = 3cos(theta)+2sin(theta) Find an expression for the slope of the curve at any point (r,Theta)

OpenStudy (stamp):

\[r(\theta)=3cos(\theta)+2sin(\theta)\]\[Find\ r'(\theta)\]

OpenStudy (anonymous):

derivative of cos(x) = -sin(x) derivative of sin(x) = cos(x) so \[r'(\theta) = -3\sin(\theta) + 2\cos(\theta)\]